### v1

# JudgeAgent: Dynamically Evaluate LLMs with Agent-as-Interviewer

Abstract
Evaluating the capabilities of large language models (LLMs) is an essential step to ensure the successful application of LLMs across various domains. The current evaluation of LLMs is based on a paradigm that involves querying them with predefined question sets and assessing their outputs. This paradigm offers controllable processes and simplicity, but faces challenges such as limited interaction with targets, insufficient difficulty control, and difficulties in verifying the validity of evaluation results, making it hard to precisely determine the knowledge and capability boundaries of target models. To address these challenges, we propose JudgeAgent, a knowledge-target adaptive dynamic evaluation framework based on a new interviewer-style evaluation paradigm. JudgeAgent employs a comprehensive evaluation approach consisting of benchmark grading, interactive extension, and evaluation feedback. It utilizes knowledge-driven data synthesis and target-adaptive difficulty adjustment methods to conduct extended testing, providing accurate and effective evaluation results. We also introduce a novel insight into validating evaluation methods, demonstrating the effectiveness of JudgeAgent and its dynamic evaluation paradigm through extensive experiments.
1 Introduction
In recent years, large language models (LLMs) have achieved remarkable success across various domains(Xu et al., 2024), from technical fields(Tang et al., 2024; Yuan et al., 2023) to social sciences(Shi, 2024; Wang et al., 2025), serving in multiple roles such as generators(Bubeck et al., 2023; Yu et al., 2024) and evaluators(Liu et al., 2023; Wang et al., 2023). This versatility stems from their massive model scale, extensive training data, and the resulting strong capabilities in comprehension and reasoning, as well as the vast amount of knowledge stored within them. Consequently, assessing the capabilities and weaknesses of LLMs has become an essential step before their deployment.
Current evaluations of LLMs are generally conducted through a paradigm that involves querying the models with predefined question sets and assessing their responses. The evaluation of outputs is typically carried out by comparing them against ground-truth reference answers, through human assessment, or via the emerging approach, LLM-as-a-judge, which uses LLMs to evaluate objects based on predefined rules and criteriaGu et al. (2024). The sources of these question sets mainly fall into two categories. The first is static datasets, which are constructed through manual collection or model-generated questions that undergo quality screening(Clark et al., 2018; Hendrycks et al., 2021; Huang et al., 2023; Lin et al., 2022; Cobbe et al., 2021; Tang & Yang, 2024). Such datasets offer advantages, including well-controlled question quality, straightforward querying procedures, and thematic consistency. However, their static nature makes them vulnerable to data leakage(Zhou et al., 2023), and easier to be cheated by target models(Zheng et al., 2024). As a result, older datasets may fail to accurately reflect the capabilities of state-of-the-art models. The second category consists of dynamically generated questions, currently also primarily produced by LLMs based on predefined requirements(Bai et al., 2023; Shi et al., 2025; Yu et al., 2024). While this approach largely mitigates the risk of data leakage, issues such as inaccuracies and irrelevance may arise due to inherent limitations in LLMs, including knowledge gaps, biases, and hallucinations(Ji et al., 2023). Furthermore, existing evaluation methods often lack sufficient interaction with the target model and do not incorporate adaptive difficulty control, making it difficult to accurately determine the model’s knowledge and capability boundaries. Moreover, the absence of interpretable feedback hinders the effective application of evaluation results to guide model iteration and optimization, and the validity of the evaluations remains challenging to verify.
To address these challenges, we propose JudgeAgent, a dynamic, knowledge-target adaptive evaluation framework with a new interviewer-style evaluation paradigm. Inspired by how a human interviewer dynamically and deeply probes a candidate’s knowledge and abilities, JudgeAgent employs a novel evaluation paradigm comprising benchmark grading, interactive extension, and evaluation feedback. By leveraging knowledge-driven synthesis of data and adaptively adjusting question difficulty based on the target model, JudgeAgent conducts extensive testing to comprehensively and profoundly assess the target model’s knowledge and ability levels. Furthermore, it provides multi-dimensional and interpretable evaluation feedback to support the target’s optimization and iteration.
In summary, our contributions are as follows:
  • Interviewer-Style Evaluation Paradigm: We propose a dynamic, strongly interactive evaluation paradigm, which comprehensively and deeply assesses target capabilities through benchmark grading, interactive extension, and evaluation feedback processes.
  • Knowledge-Target Adaptive Evaluation Framework: We introduce JudgeAgent, which dynamically adjusts difficulty based on interactive feedback and employs knowledge-driven synthetic data generation to create questions that calibrate to relevant knowledge boundaries and limits of target capabilities, enhancing evaluation validity.
  • New Insights into Validation of Evaluation: We conduct thorough experiments and analysis to validate the effectiveness of JudgeAgent’s evaluation of various target models across multiple datasets. This work also offers novel insights into how evaluation methodologies themselves can be measured.
2 Related Works
2.1 Static Benchmark-based Evaluation
These methods employ pre-constructed static benchmarks to evaluate LLMs in specific domains or on specific tasks, using formats such as multiple-choice questions, question–answer questions, or prompts that require LLMs to perform designated tasks. For multiple-choice(Clark et al., 2018; Hendrycks et al., 2021; Huang et al., 2023) or question–answer(Lin et al., 2022; Cobbe et al., 2021; Tang & Yang, 2024) formats, the benchmark provides correct answers as ground truth, and the target’s accuracy in answering the questions is used to measure its capability in the specific domain. For task-execution or open-ended formats, the benchmark typically offers only reference answers, with correctness and plausibility assessed by human evaluators(Chang et al., 2024) or LLM-based evaluators(Wang et al., 2023; Zheng et al., 2023).
The primary advantage of such methods lies in the quality of the evaluation questions, which can be well assured after prior curation. However, their static nature makes them susceptible to data leakage, whereby the target model may have prior exposure to the benchmark and can thus exploit this knowledge, undermining the validity of the evaluation.
2.2 Dynamic LLM-based Evaluation
These methods leverage LLMs to dynamically generate questions for testing target models, and in some cases, to directly serve as evaluators that provide evaluation results based on the target model’s performance. Compared with static methods, their key advantage lies in the fact that the generated questions are difficult for the target model to anticipate or tailor responses to, thereby producing evaluation results that are more likely to reflect the target’s genuine capabilities rather than its ability to recall memorized items.
For example, Bai et al. (2023) employs an LLM as an examiner, generating questions across diverse domains based on Google Trends categories, and uses follow-up questions to assess both breadth and depth of knowledge. SafetyQuizzer(Shi et al., 2025), on the other hand, retrieves up-to-date news via search engines to construct background texts and then applies retrieval-augmented generation (RAG) to formulate questions grounded in current events, maximizing question timeliness to mitigate cheating risks. These methods primarily focus on the knowledge sources used for dynamic question generation, but often lack extensive interaction and feedback with the target model. KIEval(Yu et al., 2024) emphasizes interaction with the target model by starting with known questions from a dataset and continuously issuing follow-up queries based on the target’s responses, thereby probing the depth of the target’s recall and understanding of knowledge related to the selected question. However, because it remains centered on a single question, its coverage of knowledge breadth is limited.
Moreover, dynamic evaluation methods, including the aforementioned approaches, tend to focus predominantly on the content of the questions themselves, without fully considering the alignment between question difficulty and the target model’s capability range. Evaluations conducted with difficulty levels that deviate from the target’s ability inevitably yield coarse and biased results. To address this limitation, we propose JudgeAgent, which performs authentic and fine-grained capability evaluation through iterative interaction and adaptive difficulty adjustment.
3 Methodology
Figure 1: The framework of JudgeAgent. The left part represents the interaction between the target model and JudgeAgent. The central part delineates the modular composition and operational mechanics of JudgeAgent. The right part presents the tools upon which JudgeAgent relies.
We propose JudgeAgent, a dynamic evaluation framework based on an ”interviewer-style” evaluation paradigm of benchmark grading, interactive extension, and evaluation feedback. The framework’s workflow comprises three core components: (1) Benchmark Grading: Determining the approximate capability range of the evaluation target through testing on public static datasets. (2) Interactive Extension: Using knowledge-driven synthetic methods, JudgeAgent iteratively adjusts question difficulty based on estimated target capability levels, generates extension questions for evaluation, and dynamically updates capability estimations. (3) Evaluation Feedback: Integrating all test results, JudgeAgent conducts comprehensive capability evaluations, identifies the target’s deficiencies in logical reasoning, semantic understanding, and other aspects of capabilities, and provides actionable optimization suggestions as feedback. The overall framework is illustrated in Figure X, and the detailed process is demonstrated in Algorithm 1 in Appendix B.
3.1 Benchmark Grading
Analogous to a written test before an interview, JudgeAgent, acting as the interviewer, forms an initial understanding of the target model’s capabilities by evaluating its performance on a publicly available static dataset. This process consists of two stages: Testing and Tiering.
In the Testing stage, for a selected dataset $\mathcal{D}=\{q_{1}\,,\,q_{2}\,$,\,...\}, JudgeAgent partitions a specified number of questions into batches $\mathcal{B}=\{B_{1}\,,\,B_{2}\,$,\,...\} ($B_{i}=\{q_{i1}\,,\,q_{i2}\,$,\,...\}), prompts the target model to answer the questions batch by batch, and records its responses.
In the Tiering stage, JudgeAgent determines the target model’s capability range based on its performance in different batches, thereby deciding the difficulty level on which subsequent question generation in the extension phase should be based. Within each batch, each correct answer makes the total score +1, and the final average score per question is used to evenly divide performance into three tiers: Easy, Medium, and Hard, which then guide adjustments to the question generation rules in the extension phase. In the case that the chosen dataset already provides difficulty labels for its questions (e.g., QuALITY(Pang et al., 2022)), scoring is weighted by difficulty, like Easy +1 and Hard +2, to better reflect the model’s capability level.
3.2 Interactive Extension
This process constitutes the core of JudgeAgent’s dynamic evaluation, functioning as an “interview” through which JudgeAgent conducts extensive and in-depth evaluations of the target model’s capabilities. Specifically, following the batch divisions determined in the Benchmark Grading stage, JudgeAgent iteratively expands upon each batch by generating additional questions and testing the target model. The extended evaluation proceeds in three steps: Relevant Knowledge Retrieval, Difficulty-Adaptive Question Generation, and Target Testing with Capability Estimation.
Relevant Knowledge Retrieval: Based on the original questions in a batch, JudgeAgent retrieves related knowledge from the dataset’s knowledge base. If only the reference text attached to the original questions is used as background knowledge for generating extended questions, the diversity of generated questions is constrained, leading to high similarity with the original questions and limited coverage of knowledge breadth and depth. Inspired by SoG(Jiang et al., 2025), JudgeAgent also employs a context-graph–based sampling approach for knowledge path retrieval. The context graph is constructed from the dataset’s knowledge base (e.g., the collection of all reference texts) by segmenting texts into chunks and extracting entities as nodes, with entities within the same text range as neighbors. The detailed construction process is provided in Appendix A, and is described in pseudo-code in Algorithm 3.
During knowledge-path sampling, JudgeAgent first extracts entities from the original question as seed entities, then searches the graph for the most similar entity $e$. From the text chunks associated with $e$, JudgeAgent retrieves those most relevant to the question, randomly selects a chunk $c$, and treats $(e,c)$ as the starting node of the path. The entity $e$ is then used as the new seed entity for the next sampling iteration. This process is repeated for $N$ steps to form a knowledge path with both breadth and depth. The text chunks along the path are concatenated to form the background knowledge required for question generation. In this process, the context graph ensures both knowledge breadth and depth, while the greedy similarity strategy in sampling ensures relevance to the original question. Entity extraction is performed by an LLM, with prompts detailed in Appendix C.
Difficulty-Adaptive Question Generation: Unlike other dynamic evaluation methods, JudgeAgent incorporates the target model’s capability estimation into the question-generation process to enable a more fine-grained evaluation. A fixed-difficulty test can serve a screening function by identifying underperforming models, but it offers limited utility in providing a comprehensive evaluation of a model’s knowledge proficiency. For instance, when an overly challenging test is administered to a model with limited knowledge, the evaluation may reveal that the model performs poorly and is indeed below the required competency level, but it fails to delineate the actual range of the target’s knowledge. Therefore, during both the Benchmark Grading and Interactive Extension stages, JudgeAgent dynamically adjusts the estimation of the target’s capability and the difficulty of subsequent questions based on the target’s performance across varying levels of questions. This adaptive approach allows for a precise and refined understanding of the target model’s knowledge boundaries.
Based on the capability estimation obtained from the Benchmark Grading or the previous round of Interactive Extension, JudgeAgent adjusts the difficulty of the generated questions as Easy, Medium, or Hard, and explicitly specifies the generation requirements for each difficulty level in the prompt. For Easy questions, the focus is on assessing information retrieval skills, primarily through cloze-style questions based on the original text. For Medium questions, the emphasis is on understanding concepts and key elements, with questions primarily involving basic inference. For Hard questions, the questions are designed to encourage deep thinking and complex logical analysis. Detailed prompt designs are provided in Prompt 5 in Appendix C.
Target Testing and Capability Estimation: In each round of interactive extension, JudgeAgent presents the generated questions to the target model and collects its responses. Based on responses to all questions from both the Benchmark Grading and Interactive Extension history, JudgeAgent computes scores according to the scoring scheme described in Section 3.1, taking into account both question difficulty and answer correctness. The average score is then used to determine the difficulty level for question generation in the next round of Interactive Extension.
3.3 Evaluation Feedback
This process constitutes the result–generation phase of JudgeAgent. Following the Benchmark Grading and multiple rounds of Interactive Extension, JudgeAgent aggregates all question–answer performances of the target model across varying difficulty. JudgeAgent then evaluates the target from multiple dimensions, including potential deficiencies in semantic understanding, logical reasoning, and detail capture, as well as its comprehensive performance, producing a final evaluation report in textual form. In addition, JudgeAgent simultaneously provides feedback and suggestions aimed at addressing the identified capability gaps, serving as a reference for model optimization.
Inspired by Generative Reward Model methods(Ankner et al., 2024; Ye et al., 2024; Li et al., 2025), we propose a streamlined procedure for verifying the validity of capability evaluations. For each batch of questions, JudgeAgent generates corresponding feedback and suggestions. Using these suggestions as guidance, the target model is prompted to answer the same set of questions again. By comparing accuracy before and after the intervention, we can indirectly validate the effectiveness of the evaluation to some extent.
4 Experiments
In this section, we validate the effectiveness of the evaluation results from JudgeAgent using the method introduced in Section 3.3. The detailed process is illustrated in Algorithm 2. The following research questions guide our experiments: First, does JudgeAgent genuinely discover the shortcomings of the target model (RQ1)? Second, to what extent does each evaluation mechanism in JudgeAgent influence the evaluation results (RQ2)?
Experiment Setup. We select GPT-4.1111We use gpt-4.1-2025-04-14 from OpenAI’s official API for all experiments(Achiam et al., 2023) to be the core LLM in JudgeAgent, which serves as both a generator and evaluator model. The Interactive Extension in our experiments is limited to a maximum of 3 rounds.
4.1 Dataset and Target Model Selection
The dynamic evaluation of JudgeAgent begins with Benchmark Grading, a phase that requires a predefined static dataset. In our experiments, we utilized datasets including: MedQA(Jin et al., 2021), MultiHop-RAG(Tang & Yang, 2024), and QuALITY(Pang et al., 2022).
MedQA contains multiple-choice questions in the style of the Medical Licensing Examination questions, and we select the US subset of MedQA, which has questions in English. MultiHop-RAG consists of phrase Q&A queries constructed from news articles published between September and December 2023. We use questions without reference knowledge from MedQA and MultiHop-RAG to evaluate the target model’s knowledge proficiency and to verify whether the JudgeAgent can discover the target’s knowledge gaps through dynamic evaluation. QuALITY is a multiple-choice question dataset for long document comprehension. The questions are written and validated by contributors who have read the entire passage. When employing QuALITY questions to test the target model, background passages are provided, as QuALITY is designed to assess comprehension and inference based on these passages, which primarily consist of fictional narratives. Therefore, unlike MedQA and MultiHop-RAG, we do not use QuALITY to validate the JudgeAgent’s effectiveness in evaluating knowledge gaps, but rather its ability to guide comprehension and reasoning. We select the validation set of QuALITY to conduct experiments, since the test set doesn’t provide correct answers. Although QuALITY has already categorized difficulty into 0 and 1 based on whether the annotator’s accuracy under the speed validation setting exceeds 50%, we divided the original question difficulty into three tiers of Hard, Medium, and Easy, using 20% and 60% as the thresholds to better align with the JudgeAgent setup.
For the target models, we select three LLMs as the targets: Qwen3(Yang et al., 2025), GLM4-Flash(GLM et al., 2024), GPT-4.1(Achiam et al., 2023), and gemini-2.5-pro(Comanici et al., 2025) [^2] We use qwen-plus-2025-04-28, gpt-4.1-2025-04-14, gemini-2.5-pro-preview-06-05, and the free version glm-4-flash-250414 as the target models.. We utilize the official APIs to interact with these models.
4.2 Evaluation Metrics
We utilize Accuracy to measure the performance of the target model on base datasets. To validate the effectiveness of the evaluation suggestions from JudgeAgent, we propose two metrics, Correction Rate and Correct-to-Error Rate. The detailed definition of these metrics is as follows: (1) Accuracy (ACC). This is a common metric to measure the performance, which is the proportion of questions correctly answered by the target model. (2) Correction Rate (CR). This metric quantifies the proportion of questions that the target model initially answered incorrectly but subsequently answered correctly after receiving evaluation suggestions, which measures the effectiveness of the evaluation suggestions. A higher correction rate indicates better performance of the evaluation suggestions. (3) Correct-to-Error Rate (CtE). This metric serves as the inverse of the Correction Rate, representing the proportion of questions that were initially answered correctly but subsequently answered incorrectly. A lower CtE indicates greater effectiveness of the evaluation suggestions.
4.3 Main Results and Analysis
To address RQ1, we conduct experiments on three datasets to validate the effectiveness of JudgeAgent’s evaluation, and the results are shown in Table 2 and Table 2.
Table 1: The results on MedQA and MultiHopRAG, and all values are percentages. ACC1 and ACC2 refer to the accuracy of the target model before and after receiving evaluation suggestions.
Target Model MedQA MultiHopRAG
[…]
Qwen3 91.71 96.38 5.02 0.35 63.65 70.07 17.49 11.07
GLM4-Flash 80.09 92.82 13.46 0.73 51.25 65.92 24.26 9.59
GPT-4.1 84.97 92.44 7.65 0.18 68.94 75.55 12.36 5.75
Gemini-2.5-pro 91.04 94.60 4.03 0.47 62.25 71.83 15.41 5.83
Table 2: The results on QuALITY with different difficulty levels. QuALITY-X refers to the subdataset that consists of questions with specific difficulty, and -overall refers to all questions.
Target Model QuALITY-overall QuALITY-easy
[…]
Qwen3 87.83 88.84 1.88 0.87 94.63 95.61 0.98 0.00
GLM4-Flash 73.48 76.38 4.78 1.88 83.26 84.65 4.19 2.79
GPT-4.1 89.13 92.75 3.91 0.29 93.66 96.10 2.44 0.00
Gemini-2.5-pro 93.77 96.38 3.19 0.58 97.55 99.02 1.47 0.00
Target Model QuALITY-medium QuALITY-hard
[…]
Qwen3 88.53 88.89 1.43 1.08 80.10 82.04 3.40 1.46
GLM4-Flash 74.35 79.18 5.58 0.74 62.14 64.08 4.37 2.43
GPT-4.1 91.04 94.27 3.58 0.36 82.04 87.38 5.83 0.49
Gemini-2.5-pro 93.57 95.36 2.50 0.71 90.29 95.15 5.83 0.97
Based on the results from the knowledge-intensive datasets, MedQA and MultiHopRAG, as presented in Table 2, JudgeAgent can effectively identify potential knowledge gaps in the target LLM and subsequently mitigate these gaps by providing targeted prompts that indicate possibly missing knowledge points or overlooked knowledge associations to the target models. Furthermore, as evidenced by the overall performance on QuALITY, a dataset emphasizing reasoning and comprehension, as presented in Table 2, JudgeAgent also contributes to providing further optimization guidance to address potential shortcomings in the logical reasoning and semantic understanding abilities of the target models, thereby assisting in refining the target’s thinking steps. Additionally, by comparing the Correction Rate (CR) and the Correction-to-Error Rate (CtE) of different models before and after receiving evaluation suggestions, it can be observed that the effectiveness of the suggestions is less consistent for relatively weaker models (e.g., the free model GLM4-Flash), as reflected in the higher CtE. In contrast, stronger models are less susceptible to misleading suggestions.
Based on the performance across different difficulty levels of QuALITY as shown in Table 2, we can analyze the effectiveness of JudgeAgent’s evaluation suggestions for different target models on questions of various difficulties. For relatively stronger models (Qwen3, GPT-4.1, and Gemini-2.5-pro), JudgeAgent provides stronger guidance and optimization, particularly on questions of higher difficulty (Medium and Hard). In contrast, for the relatively weaker models, JudgeAgent leads to considerable improvement across all difficulty levels, with particularly notable gains on Easy-level questions compared to stronger models. These results further indicate that JudgeAgent more accurately assesses the capability boundaries of target models and provides more difficulty-adaptive optimization to the target models. Furthermore, for all target models, JudgeAgent’s evaluation suggestions demonstrate clear optimizing guidance on Hard-level questions that involve complex reasoning, suggesting that JudgeAgent effectively identifies underlying deficiencies in target models through dynamic interactive evaluation.
In summary, the results indicate that JudgeAgent can effectively identify potential knowledge or capability gaps in target models and optimize their performance by providing targeted suggestions. Furthermore, through a difficulty-adaptive, dynamic interaction-based evaluation, JudgeAgent can more precisely delineate the target model’s capabilities, offering more refined evaluation results.
4.4 Ablation Study
To address RQ2, we conduct ablation studies on MedQA with GLM4-Flash as the target model. Under the condition of maintaining the target model’s responses to base questions unchanged, we removed different key modules and conducted the evaluation pipeline of JudgeAgent. By analyzing and comparing the performance improvement of the target model under various ablation settings, we investigated the extent to which different mechanisms influence the evaluation results. The results of the ablation study are shown in Table 3.
Table 3: The results of the ablation study with MedQA as the base dataset, and all values are percentages. $\Delta=\text{ACC1}-\text{ACC2}=\text{CR}-\text{CtE}$ refers to the overall improvement on the target models after receiving evaluation from JudgeAgent.
Target Model Evaluator MedQA
[…]
GLM4-Flash JudgeAgent 80.09 92.82 13.46 0.73 12.73
   -w/o context graph 88.60 10.26 1.75 8.50
   -w/o difficulty-adaptive 89.68 11.14 1.56 9.59
   -w/o interacitve extension 86.50 9.47 3.06 6.41
GPT-4.1 JudgeAgent 84.97 92.44 7.65 0.18 7.47
   -w/o context graph 91.36 6.72 0.33 6.39
   -w/o difficulty-adaptive 91.94 7.36 0.38 6.98
   -w/o interacitve extension 89.81 6.13 1.28 4.85
Qwen3 JudgeAgent 91.71 96.38 5.02 0.35 4.67
   -w/o context graph 95.57 4.28 0.42 3.86
   -w/o difficulty-adaptive 95.94 4.59 0.36 4.23
   -w/o interacitve extension 93.01 1.97 0.67 1.30
JudgeAgent (w/o context graph) removes the context graph and only uses chunks sampled randomly from the original knowledge base to generate questions. The context graph ensures knowledge relevance between newly synthesized questions and the base questions. The results indicate that removing the knowledge relevance provided by the context graph reduces the improvement effect of the JudgeAgent’s evaluation feedback on the target model. A possible explanation is that the context graph provides relevance between the extended questions and the base question in terms of focused entities and knowledge, thereby allowing evaluation feedback about the extended questions to supplement the background knowledge of the base question. In contrast, under this ablation setting, the lack of knowledge relevance leads to more fragmented evaluation suggestions from JudgeAgent, which may instead disrupt the thinking of target models. The higher correction-to-error rate (CtE) compared to the setting -w/o difficulty-adaptive also provides supporting evidence for this potential interference.
JudgeAgent (w/o difficulty-adaptive) removes the difficulty-adaptive mechanism and generates questions with fixed difficulty rules in the prompt. The difficulty-adaptive mechanism enables the JudgeAgent to assess the proficiency level of the target model more precisely. The results show that after removing the difficulty-adaptive mechanism, the effectiveness of JudgeAgent’s evaluation feedback was somewhat diminished. This suggests that without the difficulty-adaptive mechanism, JudgeAgent may provide evaluation feedback that deviates from the target model’s actual proficiency level, thereby compromising the effectiveness of the evaluation results.
JudgeAgent (w/o interactive extension) removes the Interactive Extension phase and only evaluates the target models with base datasets. The Interactive Extension serves as the core module of the JudgeAgent’s dynamic evaluation process, expanding both the scope of knowledge and the range of capabilities involved in the evaluation. The results show that removing the Interactive Extension significantly weakens the evaluative effectiveness of JudgeAgent. Compared to other ablation settings, a clear decrease in CR and an increase in CtE can be observed. This indicates that the comprehensive evaluation enabled by the dynamic expansion of questioning, which covers both breadth of knowledge and depth of capabilities, is crucial to the final evaluation results. Evaluating solely based on the base questions may prevent the evaluator from fully gauging the target model’s proficiency level. It is on this foundation that further improvements, such as those brought by knowledge relevance and difficulty adaptation, are built.
Additionally, the ablation experiment results comparing different models reveal that for relatively weaker models (e.g., GLM4-Flash), the performance decline after removing each component is more pronounced, which indicates that these dynamic evaluation mechanisms are particularly beneficial to weaker models and can effectively facilitate their optimization.
4.5 Case Study
To gain an in-depth understanding of JudgeAgent’s evaluation mechanism and its effectiveness, we analyze a comparative case of GLM4-Flash’s responses to MedQA questions by receiving evaluation suggestions from direct evaluation and JudgeAgent. A brief overview of the comparative case is shown in Figure 2, while the detailed content is provided in Appendix D.
Figure 2: The brief overview of the comparative case of LLM’s performance by receiving suggestions from direct evaluation and JudgeAgent.
In this case, the target LLM provided an incorrect answer when queried on a base question. When directly evaluated by an LLM evaluator without being supplied with the correct answer or relevant background knowledge, the LLM evaluator adopted a more conservative approach, in which the LLM evaluator opted to generically enumerate the potential consequences of closed abdomnial injury across various anatomical sites, rather than performing targeted differentiation between the specific medical concepts involved. Consequently, even after receiving the evaluator’s feedback, the target model continued to produce an incorrect answer. This persistence is attributable to the fact that peritonitis symptoms resulting from both colon injury and duodenal bulb injury can exhibit delayed onset, a nuance the model failed to discern.
In contrast, JudgeAgent first extracted two key entities from the base question, closed abdomnial injury and peritonitis. It then retrieved relevant knowledge entities and text chunks, such as splenic rupture and duodenal injury, from the context graph. Based on the sampled knowledge paths, JudgeAgent generated a series of extended questions. By analyzing the target model’s performance on these extended questions, JudgeAgent identified that the target model had insufficient understanding of the symptoms associated with duodenal injury caused by closed abdominal injury, thereby providing targeted knowledge guidance. Ultimately, with this feedback, the target model successfully answered the original base question correctly.
The comparative case above demonstrates that through dynamic evaluation paradigm based on knowledge-driven synthesis and difficulty-adaptive mechanism, JudgeAgent is capable of accurately identifying gaps in the target model’s knowledge and providing targeted evaluation feedback.
5 Conclusion
In this paper, we propose JudgeAgent, a knowledge-target adaptive dynamic evaluation framework based on a new interviewer-style evaluation paradigm. JudgeAgent precisely assesses the knowledge and capabilities boundaries and deficiencies of target models and provides targeted evaluation suggestions by a comprehensive evaluation pipeline, which consists of benchmark grading, interactive extension, and evaluation feedback. We also introduce a novel insight into validating the effectiveness of evaluation results. Our experiments based on these validation methods demonstrate the effectiveness of JudgeAgent and its dynamic evaluation paradigm. In our future work, we will further refine the novel evaluation paradigms and dynamic evaluation frameworks and develop more reliable evaluation tools for LLMs.
Appendix A Construction of Context Graph
For the dataset selected during the Benchmark Grading stage, the process of constructing a context graph from its knowledge base (e.g., the set of all reference texts) can be divided into three components: text chunking, node construction, and node linking.
Text Chunking: To preserve the integrity of knowledge, the granularity of chunking is not refined to the sentence level. Instead, several sentences or an entire paragraph are grouped to form a single unit, treated as the minimal hierarchical “chunk”. On this basis, chunks are further aggregated into higher-level “documents” according to whether they belong to the same article or serve as reference texts for the same question.
Node Construction: An LLM is employed to extract entities from each chunk, and the detailed prompt is provided in Appendix C. Each extracted entity serves as the subject for constructing a node. All chunks containing the same entity are assigned to the corresponding entity node. Formally, given an entity $e$, along with the chunks containing the entity $\{c_{1},c_{2},$...\} and the documents containing the entity $\{d_{1},d_{2},$...\}, the node in the context graph can be defined as follows:
$N=(e\,,\,\mathcal{C}\,,\,\mathcal{D})\ ,\ \text{where}\ \mathcal{C}=\{c_{1},c_{2}...\}\ ,\ \mathcal{D}=\{d_{1},d_{2},$...\} (1)
In practice, we store only the IDs of chunks and documents within each node to conserve space.
Node Linking: During construction, if two entities appear together in the same chunk or the same document-level texts, their corresponding nodes in the context graph are treated as neighbors. Formally, given two nodes $N_{1}=(e_{1}\,,\,\mathcal{C}_{1}\,,\,\mathcal{D}_{1})$ and $N_{2}=(e_{2}\,,\,\mathcal{C}_{2}\,,\,\mathcal{D}_{2})$, if $\exists\ c_{1i}\in\mathcal{C}_{1},c_{2j}\in\mathcal{C}_{2}\,,\,\text{s.t.}\ c_{1i}\equiv c_{2j}$ or $\exists\ d_{1i}\in\mathcal{D}_{1},d_{2j}\in\mathcal{D_{2}}\,,\,\text{s.t.}\ d_{1i}\equiv d_{2j}$, then $N_{1}$ and $N_{2}$ will be treated as neighbors in the context graph.
Appendix B Algorithm
To clarify the entire workflow of JudgeAgent, we use pseudocode to show the dynamic evaluation process in Algorithm 1, the validation process of evaluation in Algorithm 2, the construction of context graph in Algorithm 3, and the generation of extended questions in Algorithm 4.
Algorithm 1 Dynamic evaluation process of JudgeAgent
1:Input: Target LLM $\mathcal{M}_{t}$, Base dataset $\mathcal{D}=\{(q_{i},a_{i},d_{i})\}_{i=1}^{N}$ (each item include question $q$, answer $a$, and difficulty $d$), Knowledge bases $\mathcal{K}=\{k_{1},k_{2},$...\}, Core LLM of JudgeAgent $\mathcal{M}_{c}$, Predefined batch capacity $N_{B}$, Max extension round $RND_{e}$, Max hop of sampling $H$
2:Output: Evaluation Score scsc, Batched dataset with suggestions $\mathcal{D}_{S}=\{(\mathcal{B}_{1}\,,\,s_{1})\,$$\dots\}$
3:// Construct context graph
[…]
5:// Split base dataset into batches
[…]
7:// Begin Evaluation
[…]
10:for $\mathcal{B}\leftarrow\{\mathcal{B}_{1}\,,\,\mathcal{B}_{2}\,$$\dots\}$ do
11:  $sc_{\mathcal{B}}\leftarrow 0$
12:  $RND_{total}\leftarrow 0$
13:  $\mathcal{Q}_{tested}\leftarrow\{\}$
14:// Stage1: Benchmark Grading
15:  for $(q,a,d)\leftarrow\mathcal{B}$ do
16:   Get answer $a_{\mathcal{M}}\leftarrowQuery$_LLM($q\,,\,\mathcal{M}_{t}$)
17:   if $a_{\mathcal{M}}$ is correct based on $q$ and $a$ then
18:     $sc_{\mathcal{B}}\leftarrow sc_{\mathcal{B}}+$Difficulty_Score($d$)
19:   end if
20:   $RND_{total}\leftarrow RND_{total}+1$
21:   Add $(q,a,d,a_{\mathcal{M}})$ to $\mathcal{Q}_{tested}$
22:  end for
23:  Decide difficulty $d_{e}\leftarrowDecide$_Difficulty($sc_{\mathcal{B}}\,,\,RND_{total}$)
24:// Stage2: Interactive Extension
25:  for $i\leftarrow\{1,2,\dots,RND_{e}\}$ do
26:   $(q_{e},a_{e},t)\leftarrowGenerate$_Extended_Questions($\mathcal{B}\,,\,\mathcal{G}\,,\,\mathcal{M}_{c}\,,\,H\,,\,d_{e}$)
27:   Get answer $a_{\mathcal{M}}\leftarrowQuery$_LLM($q_{e}\,,\,\mathcal{M}_{t}$)
28:   if $a_{\mathcal{M}}$ is correct based on $q_{e}$ and $a_{e}$ then
29:     $sc_{\mathcal{B}}\leftarrow sc_{\mathcal{B}}+$Difficulty_Score($d_{e}$)
30:   end if
31:   $RND_{total}\leftarrow RND_{total}+1$
32:   Add $(q_{e},a_{e},d_{e},a_{\mathcal{M}})$ to $\mathcal{Q}_{tested}$
33:   Decide difficulty $d_{e}\leftarrowDecide$_Difficulty($sc_{\mathcal{B}}\,,\,RND_{total}$)
34:  end for
35:// Stage3: Evaluation Feedback
36:  Get evaluation suggestions $s\leftarrowEvaluate$($\mathcal{Q}_{tested}\,,\,\mathcal{M}_{c}$)
37:  Add $(\mathcal{B},s)$ to $\mathcal{D_{S}}$
38:  $sc\leftarrow sc\,+\,sc_{\mathcal{B}}$
39:  $N_{total}\leftarrow N_{total}+RND_{total}$
40:end for
[…]
Algorithm 2 Validation process of evaluation results from JudgeAgent
1:Input: Target LLM $\mathcal{M}_{t}$, Batched dataset with suggestions $\mathcal{D}_{S}=\{(\mathcal{B}_{1}\,,\,s_{1})\,$$\dots\},$ in which $\mathcal{B}_{i}=\{(q_{i1},a_{i1},d_{i1})\,$$\dots\}$ is a batch of base dataset $\mathcal{D}$, and $\mathcal{S}_{i}=\{s_{i1}\,$$\dots\}$ is relevant suggestions from JudgeAgent.
2:Output:Accuracy of target LLM before evaluation $acc_{1}$, Accuracy after evaluations $acc_{2}$, Correction Rate crcr, Correct-to-Error Rate cece
3:// Initialize counter of questions
[…]
5:for $(\mathcal{B},s)\leftarrow\mathcal{D}_{S}$ do
6:  $N_{\mathcal{B}}\leftarrowlen$($\mathcal{B}$)
7:  $N_{total}\leftarrow N_{total}+N_{\mathcal{B}}$
8:  for $i\leftarrow\{1,2,\dots,N_{\mathcal{B}}\}$ do
9:   $(q,a,d)\leftarrow\mathcal{B}[i]$
10:   Get answer1 $a_{1}\leftarrowQuery$_LLM($q\,,\,\mathcal{M}_{t}$)
11:   Get answer2 $a_{2}\leftarrowQuery$_LLM_with_Suggestions($q\,,\,s\,,\,\mathcal{M}_{t}$)
[…]
14:   if correct1 then
15:     $N_{acc1}\leftarrow N_{acc1}+1$
16:     if not correct2 then
17:      $N_{ce}\leftarrow N_{ce}+1$
18:     end if
19:   end if
20:   if correct2 then State $N_{acc2}\leftarrow N_{acc2}+1$
21:     if not correct1 then
22:      $N_{cr}\leftarrow N_{cr}+1$
23:     end if
24:   end if
25:  end for
26:end for
[…]
28:return $acc_{1}\ ,\ acc_{2}\ ,\ cr\ ,\ ce$
Algorithm 3 Construction process of context graph
1:Input: Knowledge Base of Dataset $\mathcal{K}=\{k_{1}\,,\,k_{2}\,$$\dots\},$ an LLM $\mathcal{M}$
2:Output: Context Graph $\mathcal{G}=(\mathcal{N},\mathcal{E})$, in which $\mathcal{N}$ is the node set, and $\mathcal{E}$ is the edge set.
[…]
4:for $k\leftarrow\mathcal{K}$ do
5:// Spliting text into chunks
6:  $C=\{c_{1},c_{2},\dots\}\leftarrowSplit$_to_Chunks($k$)
7:// Prompting LLM to label entities
8:  $P_{e}\leftarrow\emptyset$
9:  for $c\leftarrow\{c_{1},c_{2},\dots\}$ do
[…]
11:   for $e\leftarrow S_{e}$ do
12:     if $e$ not in $\mathcal{N}$ then
13:      $\mathcal{N}[e]\leftarrow(e\,,\,\mathcal{C}=\emptyset\,,\,\mathcal{D}=\emptyset)$
14:      $\mathcal{E}[e]\leftarrow\emptyset$
15:     end if
16:     Add $c$ to set $\mathcal{N}[e].\mathcal{C}$
17:     Add $k$ to set $\mathcal{N}[e].\mathcal{D}$
18:     Expand set $\mathcal{E}[e]$ with $S_{e}$
19:   end for
20:   Expand set $P_{e}$ with $S_{e}$
21:  end for
22:  for $e\leftarrow P_{e}$ do
23:   Expand set $\mathcal{E}[e]$ with $P_{e}$
24:  end for
25:end for
[…]
27:return $\mathcal{G}$
Algorithm 4 Generation process of extended questions
1:Input: Base Question $\mathcal{Q}=\{q_{1},q_{2},\dots\}$, Context Graph $\mathcal{G}=(\mathcal{N},\mathcal{E})$, LLM $\mathcal{M}$, Max hop of path $H$, Difficulty $d$
2:Output: Extended question with its answer and background text $(q_{e}\,,\,a_{e}\,,\,t)$
3:// Prompting LLM to label entities from $\mathcal{Q}$
[…]
5:for $q\leftarrow\mathcal{Q}$ do
[…]
7:  Add $set_{e}$ to $S_{e}$
8:end for
[…]
10:// Sample knowledge paths
[…]
14:for $i\leftarrow\{1,2,\dots,H-1\}$ do
15:  $E_{candidate}\leftarrowthe$ most similar 5 entities to $e^{\prime}$ in $\mathcal{E}[e^{\prime}]$ that not in $E_{visited}$
16:  $e^{\prime}\leftarrowRandom$_Sample($E_{candidate}\,,\,1$)
17:  $C_{candidate}\leftarrowthe$ most similar 5 chunks to $q$ in $\mathcal{N}[e^{\prime}].C$
18:  $c\leftarrowRandom$_Sample($C_{candidate}\,,\,1$)
19:  Concatenate $c$ to new line of $t$
20:  Add $e^{\prime}$ to $E_{visited}$
21:end for
[…]
23:return $(q_{e}\,,\,a_{e}\,,\,t)$
Appendix C Prompts
We show the detailed prompt for Entity Extraction in Prompt 4, the prompt for generating extended questions in Prompt 5, the prompt for evaluating the performance of the target LLM in Prompt 6, and the prompt for querying the target LLM with suggestions in Prompt 7.
Prompt for Entity Extraction Please identify and label the entities in the following multiple sentences, and return the entity labeling results for each sentence. The results for each sentence should be independent, in JSON format, containing the sentence number, sentence text, and the list of recognized entities (including entity text, type, and position). Return format is a dictionary, with only one key ’labeled_data’, and the value is a list, each element is a dictionary containing the sentence text and the entity list. {{   ”labeled_data”:   [     {{”text”:”Sentence 1”, ”entity_list”: [{{”entity_text”: ””, ”entity_type”: ””}}]}},     {{”text”:”Sentence 2”, ”entity_list”: [{{”entity_text”: ””, ”entity_type”: ””}}]}},     …   ] }} Notice that ”text” should be only the sentence, not the whole article. Sentence list: { Sentences }
Table 4: Prompt for entity extraction.
Prompt for Generating Extended Questions As an interviewer, you are tasked with designing questions based on the provided texts. Your role involves crafting questions and correct answers that fulfill the following criteria: 1. **Focus on the Entity**: Ensure all questions consistently center around the specified entity from the article. 2. **Ensure Accuracy and Conciseness of Answers**: Verify that the provided answer is both correct for your designed question within the context and logic of the given text fragments, and ensure the answer is sufficiently concise—presented as a word or phrase, avoiding redundancy. 3. **Conform to difficulty requirements**: You need to design questions for the required difficulty levels, with specific requirements as follows:   (1). [easy]: **Encourage Knowledge Memorization**, design questions that assess whether respondents have memorized relevant knowledge. Create questions by directly extracting and blanking out content from the given passage.   (2). [medium]: **Encourage Knowledge Comprehension**: Design questions that prompt respondents to dissect and comprehend concepts involved in the topic. Avoid assessing only superficial knowledge retention.   (3). [hard]: **Encourage Knowledge Deep Analysis**: Design questions that prompt respondents to engage in deep thinking and analysis. Avoid merely testing knowledge recall or conceptual comprehension; do not simply extract fragments from the given passage to create fill-in-the-blank items. Encourage respondents to focus on entities within the question and employ logical skills for complex reasoning. Here are examples: { Examples } Now, given the following text fragments: { context } Based on the provided texts, please generate questions by following the requirements above and referencing the examples. Output in the specified JSON format below: {{   ”generated_question”:   [     {{       ”question”: ”Generated Question”,       ”answer”: ”Correct Answer of Generated Question”     }},   …   ] }}
Table 5: Prompt for generating extended questions based on different difficulties.
Prompt for Evaluation The following is the performance of an LLM in answering a series of questions: { list of questions, correct answers, and LLM’s answers } Please evaluate and analyze the interviewee’s performance based on the above performance using concise language from the following perspectives, and provide suggestions that help the LLM answer the same questions better. Suggestions should provide specific and detailed guidance on logical thinking steps, required knowledge, and abilities, ensuring the LLM can answer correctly for the same questions. Output in the following JSON format: {{   ”flaws_knowledge”: ”The lack of background knowledge.”,   ”flaws_capability”: ”The flaws in logic and capability.”,   ”comprehensive_performance”: ”The Comprehensive performance of all questions.”,   ”suggestions”: ”Suggestions that help the LLM answer questions better.” }}
Table 6: Prompt for evaluating the performance of the target LLM.
Prompt for Querying LLM with suggestions Please complete the following question: [question]: { question } In your previous responses to these questions, the interviewer has provided the following suggestions for you to help you answer better: [suggestions]: { suggestions } Please consider the above [suggestions], and answer the above [question], in the following JSON format: {{”answer”: ”Your answer”}}
Table 7: Prompt for querying the target LLM with suggestions.
Appendix D Detailed Content of Case Study
The detailed content of the comparative case in section 4.5 is shown in Figure 3.
Figure 3: The detailed content of the case.